# Simultaneous Gesture Classification and Localization with an Automatic Gesture Annotation Model

### v1

@@ Abstract
Training a real-time gesture recognition model heavily relies on annotated data. However, manual data annotation is costly and demands substantial human effort. In order to address this challenge, we propose a novel annotation model that can automatically annotate gesture classes and identify their temporal ranges. Our ablation study demonstrates that our annotation model design surpasses the baseline in terms of both gesture classification accuracy (3-4% improvement) and localization accuracy (71-75% improvement). We believe that this annotation model has immense potential to improve the training of downstream gesture recognition models using unlabeled datasets.
Index Terms—  Gestural input, machine learning, annotation model
1 Introduction
@@
Hand gesture recognition technology is prevalent in fields like augmented reality (AR) and virtual reality (VR). Different types of hand gesture data, such as RGB, depth, and 3D skeletons, are used for recognition. With the rise of AR and VR head-mounted devices (HMDs), real-time 3D hand skeleton data is accessible and consistent across various devices and sensors [1]. In this paper, we focus on gesture classification and localization using skeleton data.
Building a gesture recognition model generally involves data collection, annotation, and training. Annotation encompasses classifying gestures and determining their temporal range. As deep-learning techniques dominate the state-of-the-art gesture recognition models [2], there’s a substantial need for labeled data. Insufficient data could lead to overfitting in deep neural networks [3]. The challenge is that data annotation, especially detailed frame-wise annotation, is time-intensive and expensive.
To combat data sparsity in gesture recognition, researchers have used data augmentation techniques like Generative Adversarial Networks for gesture data transformation [3] and unrealistic distortions like cutout [4] and mixup [5]. Approaches like few-shot learning have also been explored to reduce dataset size needs [6]. However, the reliance on existing annotated datasets and the associated annotation costs remain significant concerns. The potential of leveraging large unlabeled datasets and the automation of the annotation process remains largely untapped.
We introduce an annotation model that can predict unlabeled gesture data classification and localization. This model allows for large-scale, high-quality annotations without significant human intervention, serving as a foundation for real-time gesture recognition systems.
Using the SHREC’2021 [7] and Online DHG [8] datasets, we conducted thorough tests on our model. An ablation study revealed our model’s superior performance over the baseline model. To our knowledge, this is the first attempt at presenting an automatic model for simultaneous gesture classification and localization.
2 Related Work
@@
Traditionally, manual annotation software has been employed to annotate gestures, with human annotators manually detecting the start and end points of each gesture. However, this process is time-consuming and requires significant labor. As a result, there has been growing interest in developing automatic annotation tools. While various fields have explored this idea (e.g., [9, 10, 11]), relatively fewer works have focused on gesture data specifically. One relevant work by Kratz et al. [12] presents preliminary research on automatically segmenting motion gestures tracked by IMUs. They suggest that by recognizing gesture execution phases from motion data, it might be possible to automatically identify the start and end points of gestures. However, their work primarily focuses on gesture localization and does not address gesture classification. More closely related to our work, Lenaga et al.[13] propose the use of an Active Learning (AL)[14] framework to automatically detect sign language gesture occurrences in RGB videos. Their approach requires manual annotation for only a small subset of the videos, benefiting researchers studying multimodal communication. However, their primary application area is sign language gestures in RGB videos, which predominantly focuses on spatial data. In contrast, our model focuses on temporal data for gesture annotation.
3 Problem Formulation
@@
It is important to clarify that when referring to gesture recognition, we specifically focus on real-time gesture recognition. In this context, real-time gesture recognition involves identifying gestures within a realistic sequence that includes multiple gesture classes as well as background activities (no gestures), commonly observed in a sensor stream.
@@
Furthermore, although both gesture annotation and real-time gesture recognition aim for accurate gesture classification, they have different objectives. While real-time gesture recognition models prioritize detecting gestures as quickly as possible, gesture annotation additionally focuses on identifying the nucleus of the gesture as accurately as possible. In contrast, an annotation model can incorporate gesture nucleus localization as an inherent training objective since it does not need to optimize for recognition latency. This enables better architectural design choices to improve the performance of the annotation task.
4 Proposed Model
In this section, we describe our model for annotating gesture data.
4.1 Design Considerations
Fig. 1: Due to the extended input window of the annotation model, the true label has the capacity to encompass multiple gestures simultaneously. In the illustrated case, the true labels are {1, - , 3, - ,2}, where “-” represents background activities (no gesture). The output of the annotation model is a sequence of predicted labels that has the same length as inputs. The loss function for the proposed annotation model is a Connectionist Temporal Classification (CTC) loss. The backbone of the model is an A-ResLSTM we adopted from Shen et al. [1].
We propose an annotation model optimized for gesture classification and localization and address four key design considerations in our architecture.
@@ 4.1 Design Considerations
    Small vs Large Window Size: Conventionally, small windows ensure accurate classification by focusing on a single gesture. Although small windows lead to low recognition latency, they have limited temporal context. Hence, we utilize a larger window size with the sliding-window technique to process multiple gestures concurrently.
@@
    One Prediction vs Multiple Predictions Per Frame: Standard classification models provide a single vector of class-conditional probabilities per frame. This many-to-one mapping can lose fine-grained temporal information. Conversely, our many-to-many architecture retains richer temporal information, enabling dynamic adjustments between windows, reducing prediction noise.
@@
    Strongly-Segmented vs Weakly-Segmented Training Data: While models trained using cross-entropy loss need strongly-segmented data, obtaining such data is labor-intensive. We employ Connectionist Temporal Classification (CTC) loss, aligning unsegmented inputs with outputs without requiring strong segmentation, a technique found effective in seq2seq models.
@@
    Heuristic Parameters vs Automation: Traditional models like Shen et al. [1] employ heuristic threshold parameters to pinpoint gesture locations. As these parameters can affect localization, we favor a model devoid of heuristic parameters. Leveraging CTC loss, we directly localize the gesture nucleus by identifying the spike position in the CTC’s output without post-processing.
4.2 Annotation Model
After considering these four design factos, we now introduce our model’s individual components in detail.
4.2.1 Seq2seq Model
@@
4.2.2 Connectionist temporal classification (CTC)
Fig. 2: Demonstration of the network output probabilities from a CTC and CE (cross-entropy) trained network versus the ground truth [18]. The spike of CTC loss clearly captures the gesture nucleus, while the curve of CE loss needs post-processing. For example, the CTC loss successfully distinguishes two consecutive same gestures, while the CE loss confuses them together.
@@ 4.2.3 Dynamic Adjustment
4.2.4 Decoding
We use greedy search for the dynamic adjustment output, which approximates the solution by taking the most likely class at each step in the matrix $\mathbf{M}$. It is efficient to compute as it simply concatenates the most active outputs at each step.
5 Experiments and Analysis

### v2

# Boosting Gesture Recognition with an Automatic Gesture Annotation Framework

@@ Abstract
Training a real-time gesture recognition model heavily relies on annotated data. However, manual data annotation is costly and demands substantial human effort. In order to address this challenge, we propose a framework that can automatically annotate gesture classes and identify their temporal ranges. Our framework consists of two key components: (1) a novel annotation model that leverages the Connectionist Temporal Classification (CTC) loss, and (2) a semi-supervised learning pipeline that enables the model to improve its performance by training on its own predictions, known as pseudo labels. These high-quality pseudo labels can also be used to enhance the accuracy of other downstream gesture recognition models. To evaluate our framework, we conducted experiments using two publicly available gesture datasets. Our ablation study demonstrates that our annotation model design surpasses the baseline in terms of both gesture classification accuracy (3-4% improvement) and localization accuracy (71-75% improvement). Additionally, we illustrate that the pseudo-labeled dataset produced from the proposed framework significantly boosts the accuracy of a pre-trained downstream gesture recognition model by 11-18%. We believe that this annotation framework has immense potential to improve the training of downstream gesture recognition models using unlabeled datasets.
I INTRODUCTION
Figure 1: Our framework provides an automatic gesture annotation solution. The red borders highlight our main contribution. The framework consists of two components: (1) a novel annotation model that utilizes Connectionist Temporal Classification (CTC) loss (see Figure 2), and (2) a semi-supervised pipeline that improves the model’s performance by training on its own predictions, i.e., pseudo labels. In real-life (open-world), the gesture annotation framework operates as follows: first, a real-time gesture recognition model and our proposed annotation model are pre-trained on a small labeled dataset. Next, the annotation model is integrated into the pseudo-labeling process, where it produces pseudo labels by annotating an unlabeled dataset, which is then used to augment the annotation model. After pseudo-labeling is complete, the final high-quality pseudo-labels are used to fine-tune the pre-trained real-time gesture recognition model.
@@ I INTRODUCTION
Hand gesture recognition is a widely applied technology in various fields such as augmented reality (AR) [35], and virtual reality (VR) [39, 43]. Various modalities of hand gesture data can be utilized for recognition, such as RGB, optical flow, depth, IR, IR-disparity, and 2D/3D skeletons. Contemporary AR and VR head-mounted devices (HMDs) track the user’s hands, providing access to real-time 3D hand skeleton data. In this paper, our focus is on using 3D hand skeleton data. This choice is motivated by the widespread availability of 3D hand skeleton data in modern HMDs and its compatibility across different sensors and devices [43].
A typical pipeline to build a robust gesture recognition model usually starts with data collection, followed by data annotation and model training. The annotation step involves two steps: (1) marking the class of the gesture, and (2) localizing the time range of a gesture. State-of-the-art gesture recognition methods are primarily based on deep-learning techniques [23, 16], which require a significant amount of labeled data for training [31]. Insufficient training data can result in overfitting or a failure to learn for a deep neural-network model  [42]. Unfortunately, annotating data can be both time-consuming and costly, especially for open-world gesture recognition [46]. Additionally, many state-of-the-art gesture recognition methods require strongly-segmented data (frame-wise segmentation, i.e., Step 2 of annotation) for high-quality training [16, 4, 23], which needs more intense human manual efforts and adds another layer of expense to the data annotation process.
Previous research has focused on addressing the issue of data sparsity in gesture recognition. Various techniques for data augmentation have been proposed, such as the use of a Generative Adversarial Network (GAN) for realistic transformations on gesture skeleton data [42, 44], and the use of unrealistic distortions like cutout [6] and mixup [61] to regularize neural network training. Another recent work by Xu et al. [56] leveraged a few-shot learning framework to reduce the dataset size requirement. However, while these techniques help alleviate the issue of data sparsity, they all rely on an existing annotated dataset, which still suffers from the high cost of data annotation. There is very little research on automating the gesture annotation process that can achieve the two annotation steps simultaneously [19, 14]. Meanwhile, the efficient use of rich unlabeled datasets is underexplored.
To address the gaps, we develop a novel automatic gesture annotation framework (see the red borders in Figure 1). Our framework contains two components: 1) an annotation model (see Figure 2) that enables the prediction of unlabeled gesture data in both gesture classification (step 1 of annotation) and gesture nucleus localization (step 2 of annotation) [^1] A gesture is composed of three phases: preparation phase, nucleus, and retraction phase. Gesture localization is defined by determining the temporal position of the gesture nucleus, which is the core part of gesture recognition.. 2) a semi-supervised learning pipeline that uses the annotation model (i.e., the first component) to predict unlabeled data and then uses these predictions as pseudo-labels to further augment the annotation model itself. These high-quality labels can then be used to further fine-tune any downstream gesture recognition models (see the upper part of Figure 1).
By combining these components, we have developed a robust and automated framework for annotating gesture data of high quality. This framework enables large-scale annotation without the need for extensive human labor. It is important to emphasize that our framework primarily focuses on the annotation step, which serves as a crucial support for training downstream gesture recognition models in real-time systems.
To verify the effectiveness of our framework, we carried out an in-depth analysis using two public gesture-related skeleton-based datasets: SHREC’2021 [3] and Online DHG [5]. Initially, we conducted an ablation study to assess the contribution of our different design blocks within the annotation model on gesture classification and nucleus localization tasks. The results showed our model outperforms the baseline model by an improved gesture classification accuracy of 4.3% (3.4%) and an improved gesture nucleus localization accuracy of 71.4% (75.0%) in SHREC’2021 dataset (Online DHG dataset). Next, we used our annotation framework to label a subset of the public datasets mentioned above. We then used this pseudo-labeled dataset, created by our framework, to fine-tune a pre-trained downstream gesture recognition model. The evaluation showed that this pseudo-labeled dataset greatly enhanced the accuracy of the pre-trained downstream gesture recognition model by 11-18%.
To the best of the author’s knowledge, this is the first instance of proposing an automatic and semi-supervised gesture annotation framework that performs gesture classification and nucleus localization simultaneously.
II Related Work
We first briefly overview the related work of gesture recognition and localization. We then summarize the existing work on gesture annotation. We also introduce semi-supervised learning background.
II-A Gesture Recognition with Deep Learning
Recently, deep learning techniques have become increasingly popular for gesture recognition. Recurrent neural networks (RNNs) have demonstrated effectiveness in identifying spatial and temporal relationships in activity recognition, both with RGB-based and skeleton-based input [24, 45]. However, when there is a large gap between the input data and the target task, conventional RNNs are not capable of extracting relevant information. To address this issue, LSTMs have been proposed to manage ‘long-term dependencies’[12, 27, 41]. Additionally, graph and manifold learning has also achieved success in gesture recognition, as seen in DG-STA [4] and ST-TS-HGR-NET [32]. There has also been a growing interest in gesture recognition in AR and VR applications, such as 3D hand pose estimation using an infrared camera [33] and fast recognition of foot gestures for virtual locomotion [47].
Different modalities of hand gesture data can be used as input for gesture recognition, such as RGB, optical flow, depth, IR-left, IR-disparity, time-series data (IMU, EMG, etc.), non-optical 2D data (tomography, acoustic spectrograms, etc.), and 2D/3D skeletons [52, 59, 13, 55, 29, 26, 21, 51, 9]. The framework proposed in this paper can be easily adapted to different input types. As an illustrating example, this paper uses skeleton data, which records the 2D or 3D positions of key points/joints on the hand. With the advancement of hardware (e.g., Microsoft HoloLens, Intel RealSense, and Leap Motion Controller), the use of skeleton data is becoming increasingly popular among different platforms and modalities. These devices provide precise skeletal data of the hand and fingers in the form of a full 3D skeleton.
II-B Gesture Localization
Previous studies have explored gesture localization using RGB data, with early works such as [8, 28, 37] primarily focusing on spatial segmentation, which involves separating the hand from its surroundings. However, it is important to note that temporal segmentation, specifically determining the timing of the gesture nucleus, is equally important in addition to spatial segmentation. This temporal aspect of gesture localization is commonly referred to as “nucleus localization.” Existing approaches often employ heuristic methods to locate gestures, such as the use of sliding windows to analyze the sequence of output confidence/loss from a gesture recognition model and identify peaks or valleys [57, 58]. However, our experiments have shown that such heuristic methods face challenges when dealing with variations in gestures, as depicted in Figure 3.
II-C Gesture Annotation
@@
Traditionally, manual annotation software has been employed to annotate gestures, with human annotators manually detecting the start and end points of each gesture. However, this process is time-consuming and requires significant labor. As a result, there has been growing interest in developing automatic annotation tools. While various fields have explored this idea (e.g., [60, 34, 7]), relatively fewer works have focused on gesture data specifically. One relevant work by Kratz et al. [19] presents preliminary research on automatically segmenting motion gestures tracked by IMUs. They suggest that by recognizing gesture execution phases from motion data, it might be possible to automatically identify the start and end points of gestures. However, their work primarily focuses on gesture localization and does not address gesture classification. More closely related to our work, Lenaga et al.[14] propose the use of an Active Learning (AL) [36] framework to automatically detect sign language gesture occurrences in RGB videos. Their approach requires manual annotation for only a small subset of the videos, benefiting researchers studying multimodal communication. However, their primary application area is sign language gestures in RGB videos, which predominantly focuses on spatial data. In contrast, our framework focuses on temporal data for gesture annotation. Furthermore, our approach leverages a semi-supervised learning pipeline to enhance the performance of annotation, setting it apart from previous works.
II-D Semi-Supervised Learning Methods
In recent years, there has been a growing interest in semi-supervised Learning (SSL) due to its ability to leverage large amounts of unlabeled data. SSL becomes particularly valuable when labeled data is scarce or when the labeling process is time-consuming and labor-intensive. Consistency regularization [40, 1, 20] and pseudo-labeling [22, 54, 38] are two popular methods utilized to make effective use of unlabeled data, and they have been integrated into various modern SSL algorithms [49, 2, 53].
The semi-supervised learning pipeline within our framework draws inspiration from the recently developed FixMatch algorithm [48]. FixMatch combines these techniques with both weak and strong data augmentations, yielding remarkable results in SSL tasks.
III Problem Formulation
@@
It is important to clarify that when referring to gesture recognition, we specifically focus on real-time gesture recognition. In this context, real-time gesture recognition involves identifying gestures within a realistic sequence that includes multiple gesture classes as well as background activities (non-gestures), commonly observed in a sensor stream.
@@
Furthermore, although both gesture annotation and real-time gesture recognition aim for accurate gesture classification, they have different objectives. While real-time gesture recognition models prioritize detecting gestures as quickly as possible, gesture annotation additionally focuses on identifying the nucleus of the gesture as accurately as possible.
In contrast, an annotation model can incorporate gesture nucleus localization as an inherent training objective since it does not need to optimize for recognition latency. This enables better architectural design choices to improve the performance of the annotation task.
IV Proposed Framework
In this section, we describe our framework for annotating gesture data.
IV-A Design Questions
Figure 2: nDue to the extended input window of the annotation model, the true label has the capacity to encompass multiple gestures simultaneously. In the illustrated case, the true labels are {1, - , 3, - ,2}, where “-” represents background activities (no gesture). The output of the annotation model is a sequence of predicted labels that has the same length as inputs. The loss function for the proposed annotation model is a Connectionist Temporal Classification (CTC) loss. The backbone of the model is an A-ResLSTM we adopted from Shen et al. [43].
We propose an annotation model with individual components that are specifically designed to optimize two goals of gesture annotation: gesture classification and gesture localization. We start by answering four important model design questions of our architecture.
@@ IV-A Design Questions
    Small Window Size vs Large Window Size: The window size for a classification model is typically kept small to ensure accurate classification, as it is believed that a window should only be large enough to contain one gesture [43] such that the input data is clearly segmented and the model can be better trained. Another main reason for using a small window size is to achieve low recognition latency [30], which is not a primary concern during the data annotation process. Moreover, compared to a larger window, a small window contains less information about previous timestamps and has a narrower vision for gesture localization. Therefore, we are using a sliding-window technique with a larger window size to process multiple gestures in the input sequence.
@@
    One Prediction Per Frame vs Multiple Predictions Per Frame: When it comes to identifying gestures in a sliding-window approach, a classification model typically outputs a single vector of class-conditional probabilities for all gesture classes (plus no gesture). If the slide step is as small as one frame, the classification model can output a single vector per frame. This is a typical many-to-one mapping, as the input is a sequence of vectors, and the output is one vector. However, such a many-to-one mapping loses the fine-grained temporal information by collapsing the multiple frames into one single output. This will impact the gesture localization task. In contrast, a many-to-many (many2many) architecture of each time window contains richer temporal information and allows more dynamic and flexible adjustment between multiple windows. Therefore, we use a many-to-many model to instantiate such architecture. The model output is a sequence of class-conditional probabilities and can be decoded as a sequence of predicted classes, with overlapping predictions from consecutive windows. This allows for dynamic adjustments of earlier predictions based on later windows, and helps to reduce noise in the final prediction.
@@
    Strongly-Segmented vs Weakly-Segmented Training Data: It is important to note that a classification model that is trained using cross-entropy loss typically requires clearly defined and segmented gesture data. This means that the start and end frames of the gesture must be accurately labeled. As a result, the model requires strongly-segmented training data. However, strongly-segmented data requires extensive human labor. We thus desire a model that can be trained using weakly-segmented data, while still being able to learn the temporal and spatial properties of the gesture nucleus among the data. Therefore, to handle the challenge of the need for strongly-segmented data, we adopt a Connectionist Temporal Classification (CTC) loss, which can automatically align the unsegmented input sequence with the output sequence. This removes the requirement for strong segmentation of gestural sequences. This loss has been widely adopted in seq2seq model training [10].
@@
    Heuristic Parameters vs Automation: Shen et al. [43] used heuristic threshold parameters to estimate gesture locations. However, the accuracy of such an estimation largely depends on the other parameters of the model (window length, slide step etc.). This is because the training label does not directly contain any positional information about the gesture nucleus, and such a conventional real-time gesture recognition model is not explicitly designed for this type of localization. Therefore, we desire a model that does not need these heuristic parameters. This is another reason supporting our choice of the CTC loss. The spiky nature of the CTC’s output can eliminate the need for post-processing [50, 18], and allow for the localization of the gesture nucleus without the need for any heuristic parameters. We can identify the nucleus position directly by identifying the spike position. Moreover, greedy search can be used to decode the model output into the sequence of gesture classes.
IV-B Annotation Model
After answering these four design questions, we now introduce our model’s individual components in detail.
IV-B1 Seq2seq Model
@@
IV-B2 Connectionist temporal classification (CTC)
Figure 3: Demonstration of the network output probabilities from a CTC and CE (cross-entropy) trained network versus the ground truth [50]. The spike of CTC loss clearly captures the gesture nucleus, while the curve of CE loss needs post-processing. For example, the CTC loss successfully distinguishes two consecutive same gestures, while the CE loss confuses them together.
Connectionist temporal classification (CTC) is a type of neural network output that can be constructed into a loss function [10]. It is designed to tackle the alignment problems between two sequences with very different lengths, such as in handwriting recognition where the written text is often longer than the number of characters. This is achieved by introducing a pseudo-character (called blank --, or no label) to encode duplicate characters in handwriting recognition. In the gesture recognition case, the pseudo-character is the same as no gesture. We can then condense repeated class labels and remove no gesture labels, e.g., $\left[-,-,-,1,1,1,-,2,-,3,3,3,-,-\right]=\left[1,1,1,-,-,-,2,2,-,-,3,-\right]=[1,2,3]$, where 1, 2, 3 are actual gesture classes and “--” is no gesture.
Mathematically, we define a path $\pi$ as a possible mapping of the input sequence $\mathbf{W}$ of length $L$ into a sequence of training class labels y. The probability of observing path $\pi$ is $p(\pi\mid\mathbf{W})=\prod_{t}\mathbf{M}_{\pi_{t}}^{t},\forall\pi\in A^{\prime L}$, where $\pi_{t}$ is the class label predicted at time $t$ in path $\pi$, and $A^{\prime L}$ is the set of length $L$ sequences (paths $\pi$) over the gesture dictionary $A^{\prime}=A\cup\left\{\textit{no gesture}\right\}$. The next step is to define a many-to-one mapping operator $y=\gamma(\pi)$ to map the paths into a sequence of gesture labels, $\gamma:A^{\prime L}\mapsto A^{\leq L}$, after condensing class labels and removing no gesture labels as previously noted. Thus many paths $\pi$ under $\gamma$ result in the same gesture sequence y. The probability of observing the gesture class sequence y given an input window $\mathbf{W}$ is the sum of the condition probabilities of all path $\pi$ mapping to that sequence, $\gamma^{-1}(\textbf{y})=\left\{\pi:\gamma(\pi)=\textbf{y}\right\}$:
$p_{CTC}(\textbf{y}\mid\mathbf{W})=\sum_{\pi\in\gamma^{-1}(\textbf{y})}(p(\pi\mid\mathbf{W}))$ (1)
Therefore, the CTC loss is:
$L_{CTC}=-log(p_{CTC}(\textbf{y}\mid\mathbf{W}))$ (2)
expressed in the log domain. $p_{CTC}(\textbf{y}\mid\mathbf{W})$ can be calculated efficiently using a dynamic programming algorithm [10]. Derivatives of the loss can then be calculated for backpropagation.
@@ IV-B3 Dynamic Adjustment
IV-B4 Decoding
There are different decoding algorithms for the dynamic adjustment output. One option is greedy search, which approximates the solution by taking the most likely class at each step in the matrix $\mathbf{M}$. It is efficient to compute as it simply concatenates the most active outputs at each step. Another option is beam search. It uses more information from the output sequence by expanding all possible next steps and keeping track of the $k$ most likely steps, where $k$ is the beam factor, or the maximum number of complexes to be specialized. It requires more memory and computational power as $k$ increases. But it can generate multiple sequence candidates, which allows for selection based on other contextual information, such as a user’s posture. However, as there is currently no gesture dataset containing such context information, we leave the exploration of beam search or other decoding methods as future work and use greedy search in this paper.
Our newly proposed annotation model enables accurate gesture annotations on unlabeled datasets, including both classification and localization. These annotations can be utilized to train various downstream gesture recognition models. However, we also recognize that the potential of unlabeled datasets can be further leveraged. To address this, we introduce a semi-supervised learning pipeline that enhances the performance of both the annotation model and downstream gesture recognition models. This pipeline maximizes the benefits of unlabeled data and improves the overall performance of the framework.
IV-C Semi-Supervised Learning Pipeline with Pseudo-Labeling
Once we have an annotation model, we can apply it to unlabeled datasets and use the generated pseudo-labels to further augment the annotation model. This is a technique called pseudo-labeling in semi-supervised learning. In our case, we design a pseudo-labeling pipeline with a carefully designed training process. Specifically, we first generate two augmented versions of the same gesture data, one stronger and one weaker. We apply our seed annotation model on the weakly-augmented data and generate pseudo-labels. This is an easier task as the data is less perturbed. Then, for those labels with high confidence, we assigned them to the strongly-augmented version of the data (a harder task) and further trained the annotation model. By progressing to a more challenging task, the model can further improve its performance.
Formally, let $X=\left\{(\textbf{x}_{b},\textbf{y}_{b}:b\in(1,...,B))\right\}$ be a batch of B labeled examples, where $\textbf{x}_{b}$ are the training examples and $\textbf{y}_{b}$ are labels. Let $U={\textbf{u}_{b}:b\in(1,...,\mu B)}$ be a batch of $\mu B$ unlabeled examples where $\mu$ is a hyperparameter that determines the relative sizes of $X$ and $U$. Let $f_{seq2seq}$ denotes our proposed seq2seq model. Our approach computes a pseudo-label for each unlabeled sample which is then used in a CTC loss function. We perform two types of augmentations: strong and weak, denoted by $\mathrm{A}(\cdot)$ and $\alpha(\cdot)$.
We denote the CTC loss function (not in the negative log domain) as $p_{CTC}(\textbf{y},\textbf{x})$ from Equation 1 between label y and input x. Let $\mathbf{M}_{b}=f_{seq2seq}(\alpha(\textbf{u}_{b}))$ be the class-conditional probability distribution produced by the model given a weakly-augmented version of a given unlabeled sample: $\alpha(\textbf{u}_{b})$. Then we use $\hat{\textbf{y}}_{b}=\textit{path\_search}(\mathbf{M}_{b})$ as a pseudo-label (path_search is our greedy search algorithm in the decoding step in Sec. IV-B4). We enforce the CTC loss against the model’s output for a strongly-augmented version of the unlabeled sample: $\mathrm{A}(\textbf{u}_{b})$:
$L=\frac{1}{\mu B}\sum_{b=1}^{\mu B}\mathbb{I}(\textit{LP\_path\_search}(q_{b})\geq\tau)p_{CTC}(\hat{\textbf{y}}_{b}\mid,\mathrm{A}(\textbf{u}_{b}))$
where LP_path_search is the log-probabilities of the best path in greedy search, and $\tau$ is the threshold,
We employ basic time-series transformations [15] for data augmentation, which include scaling, shifting, time interpolation, and adding noise. These transformations help to diversify the training data and enhance the robustness of the model. For weak augmentation, we only apply noise; for strong augmentation, we apply all the transformations.
The threshold for determining if the pseudo-labels should be remained and be used to update the model controls the trade-off between the quality and the quantity of pseudo-labels. Recent work suggests that the quality of pseudo-labels is more important than the quantity for a good performance [48]. Thus we empirically set the threshold to 0.9.
V Experiments and Analysis